OpenStudy (blayne--moore):

1 over 3 y less then sign with line under it 4

OpenStudy (blayne--moore):

\[\frac{ 1 }{ 3 }y \ge 4\]

OpenStudy (aakashtomar):

Okay. Is 6 greater than 4? Yup. But what if you multiply, say, 2 to both? Even now, 12 is greater than 8. Using a similar logic, or say, property of inequalities. We can multiply 3 both sides and still the inequality will be true, right? It then becomes : \[y \ge 12\] Right? That gives you the solution set, that is, y is any number equal to or greater than 12.

OpenStudy (blayne--moore):

yes

OpenStudy (aakashtomar):

Okay, really sorry for the so many rights, was just making sure you're getting things. :P

OpenStudy (blayne--moore):

lol nah its gucci lol i understood

OpenStudy (aakashtomar):

:)